2x squared + 10x = -3
and
x squared + 10x - 7 = 0

Answer by stanbon(57347)

2x squared + 10x = -3
2x^2 + 10x + 3 = 0
Uset the quadratic formula to get:
x = [-10 +- sqrt(100-4*2*3)]/(2*2)
x = [-10 +- sqrt(76)]/4
x = [-(5/2) +- (1/2)sqrt(19)]
--------------------------------------
and
x squared + 10x - 7 = 0
x^2 + 10x -7 = 0
Use the quadratic formula to get:
x = [-10 +- sqrt(100-4*1*-7]/2
x = [-10 +- sqrt(128)]/2
x = [-5 =- 4sqrt(2)]
====================================
Cheers,
Stan H.